2^(2x)- 2^(x)- 6 = 0
Answer by khwang(438)

Rewrite 2^(2x)- 2^(x)- 6 = 0 as
(2^x)(^2)- 2^(x)- 6 = 0
Let y=2^x, then the given equation becomes
y^2 - y -6 =0,
Factor: (y-3)(y+2) = 0,
So,y =3 or -2 (since y =2^x >0, negative y is invalid.)
Hence,y = 2^x =3, and we have x = log2 3....Answer